OpenStudy (paigephay_31903):

What is the simplified expression for 3 to the power of negative 4 multiplied by 2 to the power of 3 multiplied by 3 to the power of 2 whole over 2 to the power of 4 multiplied by 3 to the power of negative 3?

OpenStudy (kenrick143):

Did you tried this?

OpenStudy (paigephay_31903):

Yes.. @kenrick143

OpenStudy (kenrick143):

Let me check

OpenStudy (kenrick143):

Did you simplify?

OpenStudy (paigephay_31903):

That's the question. It wants me to simplify. @kenrick143

OpenStudy (wckeller):

Move the 3^(-4) to the bottom, and move the 3^(-3) to the top. \[\frac{ 3^3 * 2^3 * 3^2 }{ 2^4 * 3^4 }\] Now do the operations and multiply the numbers together... \[\frac{ 27 * 8 * 9}{ 16 * 81}\] Now multiply the numbers together \[\frac{ 1944 }{ 1296 }\] This fraction in decimal form is 1.5 A.K.A \[\frac{ 3 }{ 2 }\] Final simplified answer: \[\frac{ 3 }{ 2 }\]

OpenStudy (paigephay_31903):

Thank you. @wckeller